### v1

# Exploring limits of dipolar quantum simulators with ultracold molecules

Abstract
We provide a quantitative blueprint for realizing two-dimensional quantum simulators employing ultracold dipolar molecules or magnetic atoms by studying their accuracy in predicting ground-state properties of lattice models with long-range interactions. For experimentally relevant ranges of potential depths, interaction strengths, particle fillings, and geometric configurations, we map out the agreement between the state prepared in the quantum simulator and the target lattice state. We do so by separately calculating numerically exact many-body wave functions in the continuum and single- or multi-band lattice representations, and building their many-body state overlaps. While the agreement between quantum simulator and single-band models is good for deep optical lattices with weaker interactions and low particle densities, the higher band population rapidly increases for shallow lattices, stronger interactions, and in particular above half filling. This induces drastic changes to the properties of the simulated ground state, potentially leading to false predictions. Furthermore, we show that the interplay between commensurability and interactions can lead to quasidegeneracies, rendering a faithful ground state preparation even more challenging.
Introduction — The realm of ultracold atomic and molecular systems has become a pillar for investigating fundamental quantum physics and simulating exotic many-body systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Recent advancements in controlling and manipulating magnetic atoms [12, 13, 14, 15] and dipolar molecules [16, 17, 18, 19, 20, 21, 22] have considerably extended the already very diverse array of realizable quantum systems to those that operate under long-range interactions, with remarkable achievements such as creating molecular dipolar Bose-Einstein condensates [23], dipolar supersolid phases [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], and dipolar quantum magnets [37, 38, 39]. Moreover, when coupled to optical lattices, ultracold dipolar particles can simulate long-ranged lattice Hamiltonians that mimic those encountered in condensed matter or high-energy physics that host a rich and unconventional phenomenology not observed in traditional short-range systems, but defy analytical treatment.
The tremendous technical possibilities offered by ultracold dipolar quantum simulators, though, raise the question of their accuracy. In fact, while the qualitative quantum simulation of lowest-band lattice models has been amply showcased, its quantitative correctness can be strongly affected by interactions, in particular long-ranged ones [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. It is of paramount importance to clarify this issue to capitalize on the rapid technical advancements offered by the next generation of quantum simulators and reap the rich field of long-range interacting systems. As the Bose-Hubbard model [51, 52, 53] (BHM) has been at the forefront of the quantum simulation revolution due to its fundamental applications in strongly correlated physics [54, 55, 56, 57, 58, 59, 60, 61, 62, 63], in this work we focus on the quantum simulation of its dipolar counterpart [64, 65, 66, 67, 68] that has lead to many no less groundbreaking discoveries [69, 70, 71, 72, 73, 74, 75]. We provide a systematic and quantitative comparison between the physics realized in the continuum experiment and the one observed in the corresponding lattice model. To quantify the agreement between the two descriptions, we separately calculate ground-state energies, density distributions and fidelities between many-body wave functions. This allows us to precisely delineate the experimentally relevant ranges of potential depths, interaction strengths, particle fillings, and geometric configurations where quantum simulation is accurate. Our results show that the accuracy rapidly deteriorates for shallow lattices, stronger interactions, and filling fractions above half filling. The discrepancy can be particularly high for incommensurate geometries with quasidegenerate ground states and can even lead to drastically different predictions for the density distribution. Our study provides a comprehensive chart to assess the accuracy of dipolar quantum simulators and should act as a guideline for experimental realizations. More crucially, it exemplifies the potential of such systems to realize more complex multiband lattice models hosting exotic physics.
Figure 1: Accuracy of 2D dipolar quantum simulators in different parameter regimes. (a)-(i) Many-body fidelity as a measure for the accuracy of a single-band dipolar Bose-Hubbard model quantum simulator, for varying lattice depth $A_{0}$, interaction strength $V$ and filling fraction $\nu$. The labels of (b) apply to all panels.
Physical scenario — We consider repulsive-interacting dipolar bosons of mass $m$ in a 2D optical lattice
$A(\mathbf{x})=\frac{A_{0}}{E_{R}}\left[\sin^{2}\left(\frac{\pi x}{L_{0}}+\phi_{x}\right)+\sin^{2}\left(\frac{\pi y}{L_{0}}+\phi_{y}\right)\right]$ (1)
where $L_{0}$ is the distance between two neighboring minima (unit of length) and $E_{R}$ is the recoil energy of the optical lattice (unit of energy). The optical lattice has hard-wall boundaries that restrict the number of minima to exactly $S_{x}$ in $x$ direction and $S_{y}$ in $y$ direction ($S_{i}\in\mathbb{N}$), thereby recreating a lattice with $S_{x}\times S_{y}$ sites  [^1] For an even number of sites in the $j$ direction, we need to set $\phi_{j}=\pi/2$. For the dipole-dipole interactions (DDI), we assume a strong transverse harmonic confinement which suppresses the divergence [77, 78, 79, 80] and yields the regularized form
$\mathcal{U}_{V}(\mathbf{x}-\mathbf{x}^{\prime})=\frac{VL_{0}^{3}}{E_{R}(|\mathbf{x}-\mathbf{x}^{\prime}|^{3}+\alpha)},$ (2)
where $\alpha=0.05$ and $V$ denotes the bare DDI strength between two bosons at unit distance. To limit multiple occupation of each site to $<1\%$, we add very strong contact repulsions shaped as a narrow Gaussian $\mathcal{U}_{G}(\mathbf{x},\mathbf{x}^{\prime})$ [81]. The full system is thus governed by the Hamiltonian
$\mathcal{H}_{\text{cont}}=\int\mathrm{d}\mathbf{x}~{}\hat{\Psi}^{\dagger}(\mathbf{x})\left[-\frac{\hbar^{2}}{2m}\bm{\nabla}^{2}+A(\mathbf{x})\right]\hat{\Psi}(\mathbf{x})$
$+\frac{1}{2}\int\mathrm{d}\mathbf{x}\int\mathrm{d}\mathbf{x}^{\prime}\>\hat{\Psi}^{\dagger}(\mathbf{x})\hat{\Psi}^{\dagger}(\mathbf{x}^{\prime})W(\mathbf{x},\mathbf{x}^{\prime})\hat{\Psi}(\mathbf{x}^{\prime})\hat{\Psi}(\mathbf{x}).$ (3)
Here $\hat{\Psi}^{(\dagger)}(\mathbf{x})$ creates (annihilates) a boson at position $\mathbf{x}$ and $W(\mathbf{x},\mathbf{x}^{\prime})=\mathcal{U}_{V}(\mathbf{x},\mathbf{x}^{\prime})+\mathcal{U}_{G}(\mathbf{x},\mathbf{x}^{\prime})$
To benchmark the validity of the 2D dipolar quantum simulator, we systematically map out its ground state properties in terms of three experimentally crucial parameters: 1) the depth of the optical lattice potential $A_{0}$, which we vary from $5E_{R}$ to $10E_{R}$, 2) the DDI strength $V$, which we vary in the regime of $1E_{R}$ to $10E_{R}$, and 3) different geometries (varying $S_{x}$ and $S_{y}$ independently) and number of particles, which allows us to probe a wide parameter space of filling fractions $\nu$ up to $\nu=0.67$.
Methods — To model the experimental setup, we solve the many-body Schrödinger equation directly for the continuum Hamiltonian (3) by employing the MultiConfigurational Time-Dependent Hartree method for bosons (MCTDH-B) [82, 83, 84, 85], implemented by the MCTDH-X software [86, 87, 88, 89, 90], which recasts the many-body wave function into a superposition of $M$ single-particle functions called orbitals. This allows us to obtain the continuum many-body wave function |$\ket{\Psi}_{\mathrm{C}}$ and observables derived from it, such as the total energy of the system $E$ and the particle density $\rho_{\mathrm{C}}(\mathbf{x})$ [81]. For a sufficiently deep optical lattice potential, the continuum system should map onto a BHM via a tight-binding approximation where the field operators are rewritten in terms of maximally localized Wannier functions [91, 92, 93, 94, 95, 96, 81], $\hat{\Psi}^{(\dagger)}(\textbf{x})=\sum_{\bm{\alpha}}w^{(*)}_{\bm{\alpha}}(\textbf{x})\hat{b}^{(\dagger)}_{\bm{\alpha}}$, leading to the lattice Hamiltonian
$\mathcal{H}_{\mathrm{DBH}}$ $=-\sum_{\bm{\alpha},\bm{\beta}}J_{\bm{\alpha}\bm{\beta}}\hat{b}^{\dagger}_{\bm{\alpha}}\hat{b}_{\bm{\beta}}+\sum_{\bm{\alpha},\bm{\beta},\bm{\gamma},\bm{\delta}}V_{\bm{\alpha}\bm{\beta}\bm{\gamma}\bm{\delta}}\hat{b}^{\dagger}_{\bm{\alpha}}\hat{b}^{\dagger}_{\bm{\beta}}\hat{b}_{\bm{\gamma}}\hat{b}_{\bm{\delta}}.$ (4)
Here $\hat{b}^{(\dagger)}_{\bm{\alpha}}$ denote bosonic operators for band $\sigma$ of site $j$ summarized in a unique index $\bm{\alpha}=(j,\sigma)$. The lattice model encapsulates tunneling processes between different sites and intra- or interband density-density interactions, with couplings $J_{\bm{\alpha}\bm{\beta}}$ and $V_{\bm{\alpha}\bm{\beta}\bm{\gamma}\bm{\delta}}$ obtained from Wannier function overlaps  [81].
If the quantum simulator is used to study a single-band lattice model, qualitative discrepancies can arise when the population of higher bands is non-negligible. To discern this, we consider dipolar BHM in the lowest-band approximation (1BDBH) and the “one-and-a-half band” approximation (1.5BDBH), where one band is used in one spatial direction and two bands are employed in the other. To obtain the ground state of the lattice Hamiltonian we perform exact diagonalization with the QuSpin Python library [97, 98, 99]. We then use the Wannier function basis to reconstruct continuum versions |$\ket{\Psi_{\mathrm{1/1.5BDBH}}}$ with corresponding density $\rho_{\mathrm{1/1.5BDBH}}(\mathbf{x})$. This procedure in turn allows us to compute many-body fidelities ${}_{\mathrm{C}}\innerproduct{\Psi}{\Psi}_{\mathrm{1/1.5BDBH}}$ to quantify the agreement between continuum and lattice descriptions [81].
Results — We begin by considering systems described by $M=S$ orbitals and Wannier functions (lowest-band), respectively. Fig. 1 offers an overview of the fidelity across different values of optical lattice depth, DDI strength, and filling. Each panel has been obtained from a specific lattice size and geometry ($3\times 3$, $3\times 4$, $4\times 4$ or $5\times 5$) and particle number ($N=3$ to $N=6$[81]. This variety should allow us to make general statements about the nature of the continuum-lattice mapping. The overall trend of Fig. 1 is that the fidelity decreases for shallower lattices (lower $A_{0}$) and stronger DDIs (higher $V$). The first observation is in accordance with general expectations that the tight-binding mapping is less precise for shallower lattices, since the interband gap is generally proportional to $A_{0}$. The second observation is more intriguing and is due to the long-range nature of the DDIs. As $V$ is increased, so does the interaction energy and more and more coupling terms between distant sites have to be accounted for in the lattice picture. These not only include single-band density-density interactions, but also density-assisted tunneling terms across sites and bands, which lead to a larger projection of the continuum ground state onto excited lattice bands. As a consequence, the fidelity of the continuum state with the single-band lattice state is reduced. This finding is compatible with previous observations in 1D geometries [50].
The increasing discrepancy between continuum and lattice calculations in the low $A_{0}$ / high $V$ regimes can also be evinced by comparing the energy of the corresponding ground states, shown in Fig. 2. Firstly, we find that the continuum ground state energy is systematically lower than the corresponding lattice one. This is expected since the continuum description is more fundamental than the mapping to the lattice. Moreover, as for the fidelity, the difference in energy increases with stronger interactions and shallower lattices, suggesting a less compatible description in these two cases.
Figure 2: Energetics of the dipolar quantum simulator. (a)-(i) Difference in the ground state energy between the 1BDBH model ($E_{1BDBH}$) and the continuum quantum simulator ($E_{QS}$), plotted over increasing filling fractions $\nu$.
While $A_{0}$ and $V$ play an important role in setting the accuracy of the quantum simulator, from both Fig. 1 and Fig. 2 we see that the filling fraction is the crucial quantity that mostly disrupts it. For lower filling fractions up to 0.25, the agreement is excellent with a fidelity well above 99% for most of the phase diagram. For slightly larger filling fractions up to 0.44, the fidelity remains very high in most regions and only drops to around 90% at low lattice depths ($A_{0}\lessapprox 6E_{r}$) and/or very strong interactions ($V\gtrapprox 7E_{R}$). However, filling fractions above 0.5 see a rapid decline in the fidelity to values below 50%. Interestingly, this can also occur for deep lattices with $A=10E_{R}$. The reason for this decrease should again be attributed to a sharp increase in the interaction energy – this time by forcing more particles into the same space – which grants particles the ability to overcome the band gap and excite states in higher bands, thereby reducing the lowest-band population. The dipolar quantum simulator at high densities is thus definitely breaking out of the single-band picture.
Figure 3: Improvement of the quantum simulator accuracy for the higher-band model. Many-body fidelity between continuum and 1-band or 1.5-band lattice model in a $3\times 3$ geometry as a function of $A_{0}$ and $V$ for (a) $\nu$=0.33, (b) $\nu$=0.44, (c) $\nu$=0.56, (d) $\nu$=0.67. The insets show the particle density of the single- and 1.5-band dipolar BHM for the values of $A_{0}=5E_{R}$, $V=10E_{R}$.
To quantify the contributions from higher bands, in Fig. 3 we present more detailed fidelity calculations where the 1BDBH and the 1.5BDBH are directly compared with each other for the $3\times 3$ geometry. In 1BDBH calculations, the fidelity drops monotonically for decreasing $A_{0}$ and increasing $V$ as previously seen. This worsening is particularly evident for higher filling states with $N=5$ and $N=6$, where the fidelity reaches low points of 51% and 14% respectively. However, when more bands are taken into account, the fidelity improves across all fillings. The improvement is particularly striking for $N=6$ ($\nu=0.67$). This is because the single-band lattice model predicts a drastically different ground state (with a checkerboard pattern) than the one obtained from the 1.5-band model (with a striped pattern), as can be seen in the insets of Fig. 3(d). This finding delivers two important messages. On the one hand, if the target simulation model is restricted to the single-band model, the quantum simulator can dramatically fail to obtain the correct ground states at high fillings and thus lead to wrong physical predictions. On the other hand, the physical setup is already capable of generating richer physics of more complex higher-band models without additional fine tuning.
Another key factor in the accuracy of the quantum simulator, less evident from the aggregated filling data, is system geometry. Depending on the interplay between particle number, repulsion strength, and lattice geometry, incommensurability might occur, i.e. non-integer filling for each occupied site. This can lead to quasidegenerate ground states due to spontaneous lattice symmetry breaking and thus make populating the true ground state an arduous task in an experimental realization. An overview of the quasidegeneracies emerging in the lattice system is presented in Fig. 4, where we plot the gap between the ground state and the first excited state for all the geometries considered at the four extremal values in our $(V,A_{0})$-parameter range. We can immediately distinguish two clusters of states. The clustering itself is essentially independent of the values of $V$ and $A_{0}$, thereby revealing that it is rather determined by the underlying geometry. However, within each cluster $V$ has opposite effects (stabilizing vs. destabilizing). The first cluster consists of $N=4,5$ in the $3\times 3$ and $N=4$ in the $4\times 4$ lattice, i.e. geometries that can accommodate the ground state in the most symmetric configuration in the localized limit (four isolated particles in the corners and none or one in the center, respectively). These states preserve all discrete symmetries of the lattices, such as $C_{4}$ rotations, reflections, and inversions, which makes them very stable as confirmed by a large gap to excited states. The second cluster, instead, consists of geometries with incommensurability. Examples include all the $N=3$ states and the $N=5$ state in the $4\times 4$ lattice. For these states, the localizing forces (lattice potential and repulsions) clash with the available minima compatible with the lattice symmetries. In milder cases (e.g. $N=3$, $5\times 5$ lattice), this favors a more delocalized ground state which does accommodate the symmetries but can be easily excited to similar states of quasidegenerate energies  [^2] The half-filled $S_{x}\times S_{y}=3\times 4$ case despite being commensurate was added to the incommensurate cluster due to its completely degenerate ground state.. In more severe cases (e.g. $N=6$, $3\times 3$ and $3\times 4$ lattices), a total breaking of rotational symmetries occurs, leading to a ground state partitioned into multiply-degenerate sectors, e.g. two opposite checkerboard patterns, a feature known in extended BH models for states above half filling [101, 102, 103]. Thus, we reveal another pitfall of blindly employing dipolar quantum simulators, namely the danger of preparing highly degenerate states with very different density distributions and physical properties.
[…]
Discussion and Outlook — Dipolar molecules and magnetic atoms are becoming a staple in ultracold atomic labs around the world. By transferring population to the ground state using STIRAP and controlling their large dipole moment via electric field alignment, their interaction regimes can reach $\approx 20$kHz at a $532$nm separation, i.e. $15E_{R}$ for a lattice wavelength $\lambda=1064$nm [21]. Optical lattices with hard-wall boundaries can be realized by adding flat-bottom traps to counter-propagating laser beams [104, 105, 106, 107], and lattice depths of up to $10E_{R}$ are standard. Thus, all the systems we have studied should be experimentally accessible in near-term experiments.
We have presented a comprehensive analysis of the accuracy of 2D dipolar quantum simulators by performing quantitative comparisons between continuum models of experimental setups and effective lattice models in a wide range of relevant lattice depths, interaction strengths, and filling fractions. Our study highlights that ultracold dipolar systems can be a very powerful instrument to reach regimes of multiband occupation and thereby allow to probe more realistic, complex long-range interacting lattice models. Yet, the population of higher bands can drastically alter ground-state properties, leading to erroneous predictions if the target lattice model is within a single-band picture. This is exemplified for $N=6$ particles in a $3\times 3$ geometry, where the 1.5BDBH model predicts a striped state, but the 1BDBH model gives rise to a checkerboard pattern. In general, regimes of high interaction strengths, shallow lattices, and high fillings should generate the richest multiband physics. While highly symmetric states could be simulated with greater accuracy, incommensurability effects emerging from the competition between spatial geometry and localizing interactions and potentials can drastically alter the spectral landscape. As a consequence, quasidegenerate ground states can arise, which can be completely modified by the presence of higher-band contributions. Our findings highlight a fundamental challenge in faithfully preparing and controlling single-band, long-range interacting states in ultracold quantum simulators – particularly above half-filling. This should even prompt the question whether single-band descriptions are sufficient to encompass the most exotic physics of systems (e.g. in hard condensed matter) which interact via even longer ranged couplings such as Coulomb forces. On the other hand, our results showcase the enormous possibilities offered by dipolar quantum simulators to reproduce novel physics in effective multiband lattice models.
Our study, based on a rigorous, quantitative comparison between lattice and continuum descriptions, should provide a reliable blueprint for the correct realization of long-ranged quantum simulators. While we have focused on repulsive interactions in two dimensions, we expect similar accuracy problems to arise for attractive regimes and different spatial geometries, too. This is particularly important in light of discoveries of interesting new phases in attractive dipolar BHM models such as exotic quantum liquids [75, 74, 108] and self-bound quasicrystalline order [109]. In particular, a key question to address will be the treatment of singularities arising for DDIs at zero distance, which can occur in the continuum but are truncated by the lattice formulation. Our approach of quantitatively comparing continuum and lattice descriptions by calculating many-body overlaps should shed light on those issues.
Acknowledgements.
We acknowledge computation time on the ETH Zurich Euler cluster and at the High-Performance Computing Center Stuttgart (HLRS). We thank O. Alon, E. Bergholtz, R. Chitra, D. Jaksch, and P. Nayak for useful comments on the manuscript. This work is supported by the Swedish Research Council (2018-00313) and Knut and Alice Wallenberg Foundation (KAW) via the project Dynamic Quantum Matter (2019.0068) as well as U.K. Engineering and Physical Sciences Research Council (EPSRC) Grants no. EP/P01058X/1 (QSUM) and no. EP/P009565/1 (DesOEQ).
Supplemental Material for
Exploring limits of dipolar quantum simulators with ultracold molecules
I Methods
This appendix describes in more detail the two complementary methods used to establish the validity of the dipolar quantum simulators. On the one hand, we directly simulate the continuum system consisting of optical lattice and dipolar bosons with the MultiConfigurational Time-Dependent Hartree method for indistinguishable particles. On the other hand, we diagonalize exactly the lattice model that the continuum system should map onto through a tight-binding approximation. We construct a continuum many-body wave function from the lattice ground state by expanding the lattice eigenstates with a local Wannier basis, and then compare it with the multiconfigurational wave function by evaluating the fidelity between the two many-body wave functions.
I.1 Continuum Methods
To describe the experimental setup, we solve the many-body Schrödinger equation directly for the continuum system. This is described by the following many-body Hamiltonian:
$\mathcal{H}_{\text{cont}}$ $=\int\mathrm{d}\mathbf{x}~{}\hat{\Psi}^{\dagger}(x)\left[-\frac{\hbar^{2}}{2m}\bm{\nabla}^{2}+A(\mathbf{x})\right]\hat{\Psi}(\mathbf{x})+\frac{1}{2}\int\mathrm{d}\mathbf{x}\int\mathrm{d}\mathbf{x}^{\prime}\>\hat{\Psi}^{\dagger}(\mathbf{x})\hat{\Psi}^{\dagger}(\mathbf{x}^{\prime})W(\mathbf{x},\mathbf{x}^{\prime})\hat{\Psi}(\mathbf{x}^{\prime})\hat{\Psi}(\mathbf{x}).$ (S1)
The first term describes the single-particle physics composed of kinetic energy and optical lattice $A(\mathbf{x})$, with $\mathbf{x}=(x,y)$ being the particle coordinates. The second term $W(\mathbf{x},\mathbf{x}^{\prime})=\mathcal{U}_{V}(\mathbf{x},\mathbf{x}^{\prime})+\mathcal{U}_{G}(\mathbf{x},\mathbf{x}^{\prime})$ describes the interparticle interactions, which contain both a long-range DDI repulsion $\mathcal{U}_{V}(\mathbf{x}-\mathbf{x}^{\prime})=\frac{VL_{0}^{3}}{E_{R}(|\mathbf{x}-\mathbf{x}^{\prime}|^{3}+\alpha)}$, and a short-range Gaussian repulsion $\mathcal{U}_{G}(\mathbf{x},\mathbf{x}^{\prime})=\frac{V_{G}}{E_{R}\sqrt{2\pi\sigma^{2}}}e^{-\frac{(\mathbf{x}-\mathbf{x}^{\prime})^{2}}{2\sigma^{2}L_{0}^{2}}}$ (principally needed to reduce single-site occupations – it becomes less relevant as the DDIs become stronger). We obtain the ground state |$\ket{\operatorname{\psi_{\mathrm{C}}}}$ of such many-body continuum system for various values of $N$, $A$ and $V$, and lattice geometries to compare it with the ground state obtained from the lattice model with the same parameter values.
To calculate the continuum ground state |$\ket{\operatorname{\psi_{\mathrm{C}}}}$, we employ the MultiConfigurational Time-Dependent Hartree method for bosons (MCTDH-B) implemented by the MCTDH-X software . MCTDH-X relies on a decomposition of the many-body wavefunction into an adaptive basis set of $M$ time-dependent single-particle functions called orbitals:
$|\operatorname{\psi_{\mathrm{C}}}\rangle=\sum_{\mathbf{n}}C_{\mathbf{n}}(t)\prod^{M}_{k=1}\left[\frac{(\hat{b}_{k}^{\dagger}(t))^{n_{k}}}{\sqrt{n_{k}!}}\right]|0\rangle.$ (S2)
The notation $\mathbf{n}=(n_{1},n_{2},...,n_{M})$ refers to the number of atoms in each orbital, which is subject to the global constraint $\sum_{k=1}^{M}n_{k}=N$, with $N$ the total number of particles. Additionally, the quantity $|0\rangle$ describes the vacuum and $\hat{b}_{i}^{\dagger}(t)$ denotes the time-dependent creation operator for a boson in the $i$-th working orbital $\psi_{i}(\textbf{x})$, i.e.:
$\hat{b}_{i}^{\dagger}(t)$ $=$ $\int\mathrm{d}\textbf{x}\>\psi_{i}(\textbf{x};t)\hat{\Psi}^{\dagger}(\textbf{x};t)\>$ (S3)
$\hat{\Psi}^{\dagger}(\textbf{x};t)$ $=$ $\sum_{i=1}^{M}\hat{b}^{\dagger}_{i}(t)\psi_{i}^{*}(\textbf{x};t).$ (S4)
Using the time-dependent variational principle in imaginary time, MCTDH-X optimizes both the coefficients and the orbitals by solving a coupled set of equations of motion. This procedure relaxes the system to its ground state. MCTDH-X is also able to simulate dynamics by applying the time-dependent variational principle in real-time, however this feature is not used in the present work as we focus exclusively on ground-state properties.
The MCTDH-B method allows us to calculate $N$-body reduced density matrices directly from the working orbitals. For example, the one-body reduced density matrix takes the form
$\rho^{(1)}(\textbf{x},\textbf{x}^{\prime})=\sum_{k,q=1}^{M}\rho_{kq}\psi^{*}_{q}(\textbf{x}^{\prime})\psi_{k}(\textbf{x}),$ (S5)
with
$\rho_{kq}=\begin{cases}\sum_{\mathbf{n}}|C_{\mathbf{n}}|^{2}n_{k},\quad&k=q\\ \sum_{\mathbf{n}}C_{\mathbf{n}}^{*}C_{\mathbf{n}^{k}_{q}}\sqrt{n_{k}(n_{q}+1)},\quad&k\neq q\\ \end{cases},$ (S6)
where the sum is over all possible configurations of $\mathbf{n}$ and $\mathbf{n}^{k}_{q}$ corresponds to the configuration where one particle is removed from orbital $q$ and then added to orbital $k$. The diagonal element of the one-body reduced density matrix automatically provides the one-particle density in real space:
$\rho(\mathbf{x})=\rho^{(1)}(\mathbf{x},\mathbf{x})/N$ (S7)
The number of orbitals determines the accuracy of the many-body calculations. The method is numerically exact as $M\to\infty$, but in practice a much smaller number of orbitals is sufficient to describe the correct physics. For superfluid systems exhibiting condensation, $M=1$ is equivalent to a mean-field description and is typically enough to capture the salient behaviors. For strongly localized systems in optical lattices, such as Mott insulators, it is necessary to set $M$ to be equal to the number of sites to obtain a representation equivalent to a single-band description. In our 2D calculations with $S$ sites, we employ $M=S$ or $M=2S$. This choice of $M$ thus allows us to faithfully describe each site with respectively one or two orbitals, and to project the corresponding wave function to the single-band or 1.5-band lattice basis (since a two band description of $S_{x}\times S_{y}$ sites requires $4(S_{x}\times S_{y})$ basis functions). We remark, however, that the MCTDH-B orbitals form a very different basis set than the one typically used in lattice calculations, e.g. spanned by Wannier functions. Orbitals can be highly non-local due to the variational optimization procedure in the MCTDH method. For systems that exhibit a certain degree of delocalization, as in the systems studied in this work, an orbital can incorporate more delocalized particles than, say, a maximally localized Wannier function. Thus, a lower number of orbitals $M<2S$ can be sufficient to reveal a phenomenology that appears only in higher-band descriptions for a lattice problem.
I.2 Lattice methods
To compare the physics realized in the continuum setup with the lattice models predictions, and thus evaluate the performance of the dipolar quantum simulator, we also solve the lattice model directly.
For an infinitely extended optical lattice, the single-particle Hamiltonian of Eq. (S1) possesses eigenvalues that form bands, which we can label in ascending energy order by the band index $\sigma$. Instead of employing the corresponding eigenstates, i.e. infinitely extended Bloch states, we can equivalently construct superpositions thereof that are localized at each lattice site $j$. These are known as Wannier functions $w_{j,\sigma}(\textbf{x})$.Wannier functions constructed this way from an infinite system are translationally invariant. Since in our work we deal with finite-size systems which map onto finite-size lattices, we consider a slightly modified version of Wannier functions that are built as a superposition of a particle in a box potential wave functions instead. For a detailed derivation of such finite-size Wannier functions we refer to the references  [91, 92, 93, 94, 95, 96] in the main text.
The Wannier functions allow us to decompose the bosonic field operator $\hat{\Psi}(x)$ as
$\hat{\Psi}(\textbf{x})=\sum_{j,\sigma}w_{j,\sigma}(\textbf{x})\hat{b}_{j,\sigma},$ (S8)
where $\hat{b}_{j,\sigma}$ now denotes the annihilation operator for a boson which has a spatial wavefunction given by $w_{j,\sigma}(\textbf{x})$. We can now insert this decomposition into the continuum Hamiltonian to obtain a mapping to a generalized Bose-Hubbard lattice Hamiltonian
$H=-\sum_{j,k,\sigma_{1},\sigma_{2}}J_{j,k,\sigma_{1},\sigma_{2}}\hat{b}^{\dagger}_{j,\sigma_{1}}\hat{b}_{k,\sigma_{2}}+\sum_{j,k,l,m}\sum_{\sigma_{1},\sigma_{2},\sigma_{3},\sigma_{4}}V_{i,\sigma_{1},j,\sigma_{2},k,\sigma_{3},l,\sigma_{4}}\hat{b}^{\dagger}_{j,\sigma_{1}}\hat{b}^{\dagger}_{k,\sigma_{2}}\hat{b}_{l,\sigma_{3}}\hat{b}_{m,\sigma_{4}}.$ (S9)
This Hamiltonian now consists of two processes: single-particle processes consisting of particles tunneling between different sites and bands, and two-particle processes (stemming from the continuum interactions) which take the form of intra- or interband density-density interactions (e.g. $\hat{b}_{j,\sigma}^{\dagger}\hat{b}_{k,\sigma}^{\dagger}\hat{b}_{j,\sigma}\hat{b}_{k,\sigma}$ with $j\neq k$), intra- or interband density-assisted tunneling (e.g. $\hat{b}_{j,\sigma}^{\dagger}\hat{b}_{l,\sigma}^{\dagger}\hat{b}_{j,\sigma}\hat{b}_{k,\sigma}$ with $j\neq k\neq l$), etc.
The tunnelling strength $J_{j,k,\sigma_{1},\sigma_{2}}$ between site $j$ of band $\sigma_{1}$ and site $k$ of band $\sigma_{2}$ is computed from the integral
$J_{j,k,\sigma_{1},\sigma_{2}}=-\int\mathrm{d}x\>w^{*}_{j,\sigma_{1}}(\textbf{x})\left[\frac{-\hbar^{2}}{2m}\nabla^{2}+A(\textbf{x})\right]w_{k,\sigma_{2}}(\textbf{x}).$ (S10)
The magnitude of tunnelling elements decreases rapidly as more distant sites and the bands are considered, specially for large optical lattice depths. In our work, therefore, we focus exclusively on intraband tunneling and neglect tunnelling beyond next-nearest neighboring sites. The term for $j=k$ is a particular case that corresponds to a chemical potential, which costs energy for occupation of excited bands. We remark that processes of the type where $\sigma_{1}$ not equal to $\sigma_{2}$ are zero.
The various lattice interaction terms are computed by evaluating a similar integral involving Wannier functions. For example, the lattice Hamiltonian for the DDI is calculated using
$V_{j,\sigma_{1},k,\sigma_{2},l,\sigma_{3},m,\sigma_{4}}=\frac{1}{2}\int\mathrm{d}\textbf{x}\int\mathrm{d}\textbf{x}^{\prime}~{}w_{j,\sigma_{1}}^{*}(\textbf{x})w_{k,\sigma_{2}}^{*}(\textbf{x}^{\prime})\mathcal{U}_{V}(\textbf{x}-\textbf{x}^{\prime})w_{l,\sigma_{3}}(\textbf{x}^{\prime})w_{m,\sigma_{4}}(\textbf{x}),$ (S11)
where the factor of $\frac{1}{2}$ is introduced to avoid double-counting. These elements generally have a slower decay with distance than the single-particle tunnelling elements above, due to the long-range nature of the DDI. We remark that we do not employ the common approximation that these interaction terms decay as the inverse-cube of the distance between the site minima to improve the comparison between lattice and continuum results. In our calculations, we have retained all DDI and contact interaction elements up to three consecutive sites, whereas all density-density DDI terms with $j=m$ and $k=l$ were retained since they do not depend on the spatial overlap of far-separated Wannier functions.
In the single-band model, these terms correspond to density-density repulsion $\hat{n}_{i}\hat{n}_{j}$, while for higher bands, extra terms appear describing bosons tunneling between the two bands within a site.
When all the bands are taken into account, the lattice description is completely equivalent to the continuum one. However, this limit is computationally intractable. By restricting the number of bands to only a few, we end up with a (at least numerically) solvable effective lattice models that describe the physics of long-range interacting particles tunneling across the lattice. These are the models routinely employed to study the phenomenology of numerous condensed matter systems and try to understand the microscopic origin of macroscopic phenomena such as superconductivity, magnetism etc. In the very deep lattice limit, where the potential energy dominates over all other energy scales (including interactions and thermal fluctuations), accounting for a single band only is typically sufficient to give a qualitatively correct description of the lattice physics. When the interaction energy becomes more relevant (either directly by increasing the interaction strength and the involvement of the long-range tail of the DDI, or indirectly by considering shallower lattice and/or higher particle densities), particles can hop to higher bands and these have to be considered. If the dipolar quantum simulator is thus employed to study a single-band lattice model, qualitative discrepancies can (and typically will) arise. To discern between these two cases and map out the validity of the quantum simulator, we have solved the lattice model within a single-band approximation and what we call the one-and-a-half band approximation, where one band is used for one spatial direction and two bands are employed in the other. In our simulations, we employed 2 bands per site in $x$ and 1 band per site in $y$ direction. However, we remark that the results should be equivalent for the opposite choice (2 bands per site in $y$ and 1 band per site in $x$ direction) since most of the systems we considered are symmetric with respect to a coordinate swap $x\leftrightarrow y$. The only system which is affected by the choice of the direction for the two bands is 6 particles in a $3\times 3$ geometry. As the ground state for this system are stripes, the choice of the direction of the two bands changes the orientation of the stripes. The choice of the number of bands results in two separate kinds of dipolar Bose-Hubbard models which we label as lowest-band dipolar BH model (1BDBH) and lowest-1.5-bands dipolar BH model (1.5BDBH), respectively.
To solve the lattice models we employ exact diagonalization (ED) performed numerically by the QuSpin Python library. As the repulsive DDI and contact interaction strongly discourage multiple occupation of sites, we limited the number of bosons in each site of each band to $2$. In the main text, the resulting wavefunction is denoted by |$\ket{\operatorname{\psi_{\mathrm{ED}_{1}}}}$ for the 1BDBH and by |$\ket{\operatorname{\psi_{\mathrm{ED}_{1.5}}}}$ for the 1.5BDBH.
We finally add a technical remark for the reason why we accounted for higher bands in one direction only. To obtain a 1:1 correspondence between the continuum and lattice model, we necessitate the same number of basis states (orbitals in the continuum and Wannier functions in the lattice) in both settings (see also next section). For the number of sites considered in this work ($S=9$ to $S=25$), this requires $M=36$ to $M=100$ basis functions. The computational complexity of MCTDH-X is determined by the many-body configurations, which are given by $\left(\begin{array}[]{c}N+M-1\\ N\end{array}\right)$ and the computational cost scales as $M^{4}$$\left(\begin{array}[]{c}N+M-1\\ N\end{array}\right)$. For $N=3$ to $N=6$ particles considered, this scaling limits the number of orbitals to around 20 to 25 (depending on $N$). Since this number has to correspond to the number of Wannier states, with current state-of-the-art computational resources even for $S=9$ the access to the full 2-band picture (requiring $M=36$) is precluded.
I.3 Method comparison
We employ many different observables to quantify the agreement between lattice and continuum calculations. The most immediate quantity is given by the ground-state energy, which is given by the sum of kinetic, potential, and interactions energies. Since the continuum model is the more fundamental description of the experimental system that we aim to reproduce, we find that the energy of the MCTDH-X calculations is systematically lower than the one obtained from the ED of the lattice model. However, the quantitative discrepancy is in general quite small when the lattice and the continuum model agree with each other, and increases as the agreement is progressively lost.
The second quantity we use to compare the two models is the real-space bosonic density. In MCTDH-X, this is calculated directly from Eq. (S7). In the lattice calculations, we first obtain the lattice wavefunction
$|\operatorname{\psi_{\mathrm{ED}}}\rangle=\sum_{\mathbf{n^{\prime}}}C^{\prime}_{\mathbf{n^{\prime}}}\prod^{L}_{k^{\prime}=1}\prod^{N^{\prime}_{b}}_{\sigma^{\prime}=1}\left[\frac{(\hat{b^{\prime}}_{k^{\prime},\sigma^{\prime}}^{\dagger})^{n^{\prime}_{k^{\prime},\sigma^{\prime}}}}{\sqrt{n^{\prime}_{k^{\prime},\sigma^{\prime}}!}}\right]|0\rangle$ (S12)
from ED. In Eq.S12, we have used primes to distinguish lattice quantities, i.e. $N^{\prime}_{b}$ quantifies the number of bands in the lattice model, $\mathbf{n^{\prime}}=(n^{\prime}_{1,1},...,n^{\prime}_{k^{\prime},\sigma^{\prime}})$ is the number of bosons in each Wannier function (subject to the global constraint $\sum_{k^{\prime}=1}^{L}\sum_{\sigma^{\prime}=1}^{N_{b}}n^{\prime}_{k^{\prime},\sigma^{\prime}}=N$, with $N$ the total number of particles), and $\hat{b^{\prime}}_{i^{\prime},\sigma^{\prime}}^{\dagger}$ denotes a bosonic creation operator for the $i^{\prime}$-th Wannier function of the $\sigma^{\prime}$-th band $w_{i^{\prime},\sigma^{\prime}}(x)$. Then, we construct $\langle\operatorname{\psi_{\mathrm{ED}}}|\hat{\Psi}^{\dagger}(x)\hat{\Psi}(x)|\operatorname{\psi_{\mathrm{ED}}}\rangle$ by employing the field operator from (S8) evaluated on the same spatial grid used by the MCTDH-X calculations. This gives a quantitative measure to compare the spatial distribution of the ground state across both models and provides an immediate, visual comparison for potential discrepancies.
Finally, to compare the similarity of the full many-body wavefunction across both models, we construct many-body fidelities. This is achieved by first projecting the MCTDH-X wavefunction |$\ket{\operatorname{\psi_{\mathrm{C}}}}$ to the Hilbert space of the lowest band to obtain |$\ket{\operatorname{\psi_{\mathrm{C}\rightarrow\mathrm{ED_{1}}}}}$ or of the lowest 1.5 bands to obtain |$\ket{\operatorname{\psi_{\mathrm{C}\rightarrow\mathrm{ED_{1.5}}}}}$. We then compute the fidelities of these projections with the ED wavefunctions as $f_{1}=|\langle\operatorname{\psi_{\mathrm{ED}_{1}}}|\operatorname{\psi_{\mathrm{C}\rightarrow\mathrm{ED_{1}}}}\rangle|^{2}$ and $f_{1.5}=|\langle\operatorname{\psi_{\mathrm{ED}_{1.5}}}|\operatorname{\psi_{\mathrm{C}\rightarrow\mathrm{ED_{1.5}}}}\rangle|^{2}$. The inner product used in these fidelity calculations is the one for the single-particle basis functions for the two methods i.e. the working orbitals $\psi_{i}(x)$ for the continuum calculations and the Wannier functions $w_{i,\sigma}(x)$ for the lattice calculations. This can be obtained from the overlap integrals
$O_{i,(i^{\prime},\sigma^{\prime})}=\int\mathrm{d}x\>\psi_{i}(\textbf{x})^{*}w_{i^{\prime},\sigma^{\prime}}(\textbf{x}),$ (S13)
which are the coefficients of the projection of the annihilation operator for the MCTDH-X orbitals into the lattice Hilbert space
$\hat{b}_{i}=\sum_{i^{\prime},\sigma^{\prime}}O_{i,(i^{\prime},\sigma^{\prime})}\hat{b}_{i^{\prime},\sigma^{\prime}}.$ (S14)
With this formula, the projection of the MCTDH-X wavefunction onto the lattice Hilbert space |$\ket{\operatorname{\psi_{\mathrm{C}\rightarrow\mathrm{ED}}}}$ can be computed from equation S2. The fidelities then follow from a lattice Hilbert space calculation.
II Many-body fidelities
In this section, we present a comprehensive review of all the fidelities between the continuum and lattice states calculated to generate the phase diagrams in the main text. The range of different filling fractions was achieved by varying the geometry and the particle number as shown in table S1.
Table S1: Geometries and filling fractions
Geometry Number Number of Filling
($S_{x}\times S_{y}$) of sites particles (N) fraction ($\nu$)
5 $\times$ 5 25 3 0.12
4 $\times$ 4 16 3 0.19
4 $\times$ 4 16 4 0.25
4 $\times$ 4 16 5 0.31
3 $\times$ 3 9 3 0.33
3 $\times$ 3 9 4 0.44
3 $\times$ 4 12 6 0.5
3 $\times$ 3 9 5 0.56
3 $\times$ 3 9 6 0.67
Figure S1: Subplots (a)-(i) show how the lattice depth ($A_{0}$), dipolar interaction strength ($V$) and the filling fraction ($\nu$) affect the fidelity. The interplay of high $V$, low $A_{0}$ and high $\nu$ are found to be able to reduce the efficacy of a quantum simulator.
As we can see from the plots, the fidelity tends to decrease when simulating systems with stronger interactions or in shallower lattices. This trend is rather systematic and can be observed for almost all cases. The change in curvature observed in certain plots (e.g. very low filling fractions for $N=3$ particles in a $5\times 5$ geometry) can be attributed to numerical imprecision and approximation of the true continuum with a limited number of bands, that manifests itself when the fidelity takes already exceptionally high values above 99.9%. The calculations for the half-filled lattice points $A_{0}=5E_{R}$, $V=1E_{R}$; $A_{0}=5E_{R}$, $V=2E_{R}$ and $A_{0}=5E_{R}$ and $V=10E_{R}$ have produced zero fidelity values. We conjecture that the first two points did not give meaningful results due to numerical effects in the shallow lattice regime for this degenerate ground state. Therefore for the fidelity plots in the main paper we have chosen to interpolate these two points. The last point is zero due to the chosen geometry (3 $\times$ 4 lattice), which for the checkerboard ground state configuration has two empty corner sites. As $V$ increases, the orbitals in the continuum calculations (which are more delocalized due to the optimization procedure) are pushed closer to the borders while the lowest-band Wannier functions, being site-centered in nature, do not exhibit this behavior. At the critical point of $V=10E_{R}$ the checkerboard is destroyed, as occupying the empty corner site is preferred in the continuum calculation. We however believe that for larger lattices that might be used in experiments this effect will not play a major role.
III Real-space density distributions
In this section, we present a visual depiction of the accuracy of the quantum simulator by directly comparing the real-space density distributions of the ground states obtained from the continuum with the ones calculated from the lattice model. For each geometry and particle filling, we visualize both the best and the worst scenario by comparing the density distributions side by side. Fig. S2 shows the single-band while Fig. S3 shows the 1.5-band visual comparisons. The most striking difference between the states is the case of the filling fraction $\nu$=0.67 for the single-band model, where the 1BDBH model is a checkerboard-like pattern, while the continuous quantum simulator calculation produces stripes. Increasing number of bands gives the stripe pattern for the 1.5BDBH model which is shown in Fig. S3. The discrepancy for the half-filled case for the $A_{0}=5E_{R}$ and $V=10E_{R}$ values in Fig. S2 is due to the rectangular geometry of the lattice as was discussed in the previous section.
Figure S2: Probability density of the exactly solved single-band Dipolar Bose-Hubbard model (columns 1 and 3) and the continuum calculation representing the quantum simulator (columns 2 and 4) for parameter regimes $A_{0}=10E_{R}$, $V=1E_{R}$ (best fidelity) and $A_{0}=5E_{R}$, $V=10E_{R}$ (worst fidelity). The rows are plotted in increasing values of filling fraction ($\nu$).
Figure S3: Probability density of the exactly solved 1.5-band Dipolar Bose-Hubbard model (columns 1 and 3) and the continuum calculation representing the quantum simulator (columns 2 and 4) for parameter regimes $A_{0}=10E_{R}$, $V=1E_{R}$ (best fidelity) and $A_{0}=5E_{R}$, $V=10E_{R}$ (worst fidelity). The rows are plotted in increasing values of filling fraction ($\nu$).
IV Orbital convergence
In this section, we present continuum results that illustrate how several quantities converge as a function of number of orbitals. For simplicity, we restricted our calculations to the point in parameter space having the worst many-body fidelity with the single-band lattice wave function, namely $N=6$ particles in a $3\times 3$ geometry (corresponding to filling fraction $\nu=0.67$) with $A_{0}=5E_{R}$ and $V=10E_{R}$. Another reason for considering this point is the fact that we can rather easily perform MCTDH-X calculations up to $N=6$ particles with $M=22$ orbitals, and thus probe a larger convergence regime, across single-band and 1.5-band calculations in equivalent lattice formulations and beyond. This is not possible for larger system sizes (e.g. $5\times 5$ sites) where probing a number of orbitals beyond an equivalent number of single-band Wannier functions ($M=25$) is numerically prohibitive (and 1.5-band calculations in the lattice picture are also beyond current numerical capabilities).
Fig. S4 shows the convergence results for the total energy [panel (a)], the orbital occupation [panel (b)], and real-space density distribution [panel (c)]. As we can see in Fig. S4(a), the ground-state energy converges exponentially in the number of orbitals. The density in Fig. S4(c) is initially different than the striped configuration obtained for the 1.5-band lattice density observed in the main text. In fact, it seems that the system is trying to interpolate between the checkerboard pattern (with on average one particle in the middle, four in the corners, and another particle delocalized across the remaining sites) and the striped pattern (three particles along opposite edges). However, with increasing $M$, the density distribution rapidly crystallizes into its striped configuration. This happens already at $M=8$ (with less than one orbital per site, i.e. an even more succinct description than the single-band lattice picture), and retains the same structure all throughout $M=22$. This is remarkable, given that we still observe a sizeable decrease in energy in the same regime, and indicates that MCTDH-X should be an excellent instrument to investigate ground-state configurations even with a number of orbitals far lower than the number of bands typically needed in the corresponding lattice models. This striking convergence is also observed in the orbital occupation [Fig. S4(b)]. Due to the strong particle localization, most (around 99.8%) of the many-body state is described by the most occupied 6 orbitals already at $M=10$. The more orbitals are introduced, the smaller their contributions become. For instance, from the 17th orbital onwards, all additional orbitals contribute to less than 0.05% of the total many-body state, indicating a dramatic orbital convergence of the continuum numerics.
Figure S4: Convergence of MCTDH-X calculations for $N=6$, $A_{0}=5E_{R}$, $V=10E_{R}$ as a function of number of orbitals $M$ for several observables: (a) total ground-state energy, (b) orbital occupation, (c) real-space density distribution.